Let W be the subspace spanned by the given vectors. Find a basis for Wt, 0 1 A. W1 = W2 = 3 2 -1 2 B. W W2 2 -3 W3 = 6
Answer #1

Given wie and W₂ = 3 - 2 ( .2 we know that lasta {well; <w.u3=0,*4EW} so we see all vectors (x,y,z, *) such that <w,w,>co and(1) Set 2:1, 6:0 from yozz+tzo 9 -20 5) y: 2 also x-y + 32-24:0 X-2 +3 -0 =0 X+1=0 X : (2,2,1, 0) is Solution 1:4 (ii) Se+ 2:{-100°171-7 PM २५+ .5B wi- [1] -- [1-6] وله we seek all vectors (x,y, 2, k) such that (2,4, 2, 6)· (2,-), 6,3) 20 (2,4,2,4). (-1,2,-3,-2) =0 (2,4,1) C 2:0 from 39 - - ㅋ 39-150 방할 from 21 y + 62 43t:0 9 2% x to +320 ) 2X+ +34 2 + . 키 22 키 20 - XE : (1) is o Solution u - {